Home‎ > ‎Reports‎ > ‎Crew 109‎ > ‎

day03


Crew #109 Log Book for December 19, 2011

Photos of the day





Commander's Check-In Report

Date and time: 19 December 201, 17:10

Written by: Nicky De Munster, Crew 109, Commander

Crew Physical Status: Good

Time Departed/Returned from EVA:

EVA 1: 8:45 - 11:10 am
EVA 2: 1:10 - 4:30 pm

Brief Narrative of Field Mission Results:

Both EVA teams explored the environments and took rock and soil samples at different places. The soil will be used for seeds to plant.

David will investigate the soil samples for water flow. The places were recorded by the GPS system.

EVA Data/Interpretations:

Soil will be used for seeds to plant tomorrow
Rocks samples with penetrometer on two different spots

Engineering/Hab Maintenance:

Water trailer empty
ATV's: fuel jerrycans empty

Report Transmission Schedule:

Plans for Tomorrow: 2 EVA's; lab: seeds, water analyses

Inventory: n/a

Support Requested:

Water trailer is empty
Key for lab cabinet 8 (microscopes)
Low on ATV fuel, only for 2 EVA's left
Kitchen: can opener broken

Miscellaneous:

Commander's Report

Date and time: 19 December 2011

Written by Nicky De Munster, Crew 109, Commander:

We had two great EVA's today. The two crews drove to Sagan Street and on some places they found muscovite. We called this place Cristal Field.
We find that we can do no more than two EVAs each day. It is very exhausting and there is also the problem with the fog on the helmets.
As long as the weather is good, we go on EVA.
Analysing samples will start tomorrow.
Yesterday we started the food study. The food is good and there is plenty of choice.
Karon made ​​a good daily schedule with crew rotations for EVA's, cooking and cleaning.
The mood is very good and we are all excited about the environment.

Science Report

Date and Time: 19 December 2011, 19:30
Written by David Kutai Weiss, Crew Geologist & Executive Officer, Karon Wynne, Health & and Safety Office and Crew Biologist, Victoria Grudzinski, Crew Biologist

Biology Report:
On both EVA missions today the Biology team collected more soil and water samples for our studies. We are in the process of getting plates ready for plating our soil samples tomorrow. The team also made a very exciting discovery, an entire field of fossilized bivalves. This is indicative of a great body of water in the area at some point in history. Victoria and Karon will attempt to identify the species of the specimens. Karon collect a soil sample at the location and 3 mollusk fossils at another location.


Geology Report
The second EVA of the day set out with the goal of collecting 2 core samples and 6 soil strength measurements. This was a lofty goal, as the process of collecting samples is physically taxing and the 2 locations for measurements and collections were miles apart. The crew headed straight to the first location, Site A where they conducted 3 trials of the blowcount test with the cone penetrometer and took a core sample. The soil samples were each put into 2 ziplock bags which were labeled. With the day light hours dwindling the crew hurried to the second collection site at Site F. After repeating the same procedure from the first site, the crew made it safely back to the Habitat just as the sun was going over the hill.
These samples will be used to determine water content of their respective soils. The penetrometer data was entered into an excel spreadsheet. Ultimately, this data will be used to create a soil mechanics regression chart from which soil strength can be inferred. Vast areas of bivalve fossils were observed along Radio Ridge Road. We also traveled very far North along Skyline Rim to a patch of very large and pristine muscovite mica samples exposed in the surface soil. David brought a very large piece of muscovite mica to the Hab.


EVA 3 & 4 Report

Date and Time: 19 December 2011, 19:00
Written by David Kutai Weiss, Crew 109 Geologist & Executive Officer

EVA 3:
Time: 8:42 am – 11:00 am
Crew Members: Nicky, Aster, and Victoria
Site Location: West of Hab at the edge of Skyline Rim
Transit Mode: ATV, on foot.
Objectives: Soil sample acquisition to plant seeds, search for water.

Results:

Route spanned 13 km.
Nicky took soil samples, found pristine muscovite mica samples, bivalve fossils.
Victoria took a snow sample, a soil sample, and fossil sample.
Aster had a successful test of experimental EVA helmet camera.

EVA 4:
Time: 1:15 pm – 4:30
Crew Members: David, Leandro, and Karon
Site Location: Traveled south on Radio Ridge Road toward soil sample site B. Traveled to Skyline Rim to sample site F, and as far North on Copernicus Highway to a patch of very pristine muscovite mica crystals.
Transit Mode: ATV, on foot.
Objectives: General North-western exploration and cresting Skyline Rim. Travel to site A and Site F to perform a blowcount test with a cone penetrometer and to acquire core soil samples.

Results:

Route spanned 22 km, we collected soil core samples and performed 3 trials of the blowcount test with the cone penetrometer at both site A and F. We also traveled very far North along Skyline Rim to a patch of very large and pristine muscovite mica samples exposed in the surface soil.
Karon took soil sample at Site A, and collected a bivalve sample at site F.
Leandro collected bivalve samples at Site F, and observed deer and some sort of canine (coyote?) footprints along Radio Ridge Road.

EVA 3 & 4 Lessons learned: 

Plan EVA routes in advance for a more efficient transit and sample acquisition. Leave ATV in gear when stopped, or engage handbrake when leaving ATV – even on seemingly flat areas, the ATV may move. Do not get off the ATV while it is running. Bring a camelback or some type of water bladders for EVA’s. Long EVA’s require backup batteries for the radios.

Engineering Report

Date and time: 19 December 2011, 18:33

Written by Leandro G. Barajas, Crew 109, Chief Engineer:

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): 0.25
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): Empty

Notes/Comments:
All 3 ATVs fuel tanks are full for tomorrow's EVA however there is no gas left in any of the jerrycans or the main gas tank.
Warning: We will need to get resupplied with gas by tomorrow or we will have to stop ATV EVAs after that.


Oil:
ATV Oil Quantity (# of Quarts in storage): 1.3
Car Oil Quantity (# of Quarts in storage): N/A

Water Status:
Outside Potable Water Tank Level (inches from bottom): 15
Trailer Potable Water Tank Level (inches from bottom): 4 (it is not possible to pump more water out of it)
Hab Potable Water Tank Level (inches from bottom): 11
Potable Water Meter Reading: 478577
In to GreenHab Meter Reading: N/A

Notes/Comments:
CRITICAL: We need to get resupplied with water in the next 24-48 hours.

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A
Greenhab Notes/Comments: N/A

Observatory:
Telescope Used? (Yes/No): NO

Observatory Notes/Comments: N/A

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): YES (moved a few feet away from Hab)
Oil Checked (yes/no): NO
Tire Status: Good
Odometer at end of day: N/A
Notes/Comments on Hab Car: N/A

ATV 1 (Opportunity):
Used (yes/no): YES
Oil Checked (yes/no): YES
Fuel Consumed: 1.3 Gal
Tire Status: Good
Comments:

ATV 2 (Spirit):
Used (yes/no): YES
Oil Checked (yes/no): YES
Fuel Consumed: 1.3 Gal
Tire Status: Good
Comments:

ATV 3 (Viking I):
Used (yes/no): YES
Oil Checked (yes/no): YES
Fuel Consumed: 1.3 Gal
Tire Status: Good
Comments:

Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 73

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 63

Computers/Networking Infrastructure:

Notes/comments on internet/computers?:
- During day 1 after we connected all computers, we think one of them got some kind of automatic update and the internet got pretty slow after that.
After last night, the internet speed was normalized.

General Engineering Notes/Comments:

Heat Tapes Operating 24/7 (Both Lights On)
Troubleshooted radios and headsets today. These are the results:
- 2 Radios have problems and do not work with rechargable batteries packs. However they are operational using non-rechargable batteries.
- 2 Radios have intermitent Tx/Rx problems
- 2 Headsets are dead
- 2 Headsets only the speaker works, microphone is dead.
- 4 Headsets work only when VOX mode is selected in the radios.No headsets can be used with manual PPT.

Journalist Report

19 December 2011, 18:34

Written by: Aster Stein

Today was the first full day under the food study regimen. Microwaved rehydrated whole eggs with bacon were just what we needed as fuel for what would prove to be an action packed day. Today we had the same EVA-crew as yesterday, but on the other time slot. Nicky, Victoria and Aster kicked off the AM EVA-shift, aiming for the very end of a track called Sagan Street. The tracks were very hard to see, so initially it took some time to get going. En route, they found some very interesting area's, henceforth known as "Crystal Field" and "Fossil Heap".

During the EVA-crew's absence, the Hab-crew made a delightful culinary effort consisting of tortilla's with nutella, peanut butter, left-over chili and left-over brownies. A hearty lunch all in all. The PM EVA-crew had an exciting trip as well. They dragged David's geology equipment out of the airlock and loaded it on the ATV's and gathered wonderful data for his mapping project. They also discovered a field of giant muscovite slabs and dubbed it "Mirror Field".

Before allowing themselves to eat a well earned diner, the entire crew had a lot of homework to turn in. Reports were written, thoughts about today's activities were shared. That's when gems like this spontaneous philosophical self-reflection in haiku form(by David) were born:

Today soil mapping
Tomorrow more soil mapping
David quite content.

Other